OpenStudy (anonymous):

The hypotenuse of a right triangle is 25cm, and the shorter leg is 15cm. Find the length of the remaining leg. A. 2 (Sqrt)5 cm B. (Sqrt)10 cm C. 20 cm An explanation how to reach the answer, and the answer would be appreciated. Medal given of course.

OpenStudy (help!!!!):

basically use a^2+b^2=c^2

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm a^2+b^2=c^2 }\) \(\large\color{midnightblue}{ \rm 15^2+b^2=25^2 ~~⇒~~\sqrt{25^2-15^2} ~~⇒~~?? }\) you tell me

OpenStudy (anonymous):

So, 25^2 + 15^2 = the last leg?

OpenStudy (anonymous):

Oh okay. I see.

OpenStudy (anonymous):

So, the result of that is 400. And when you find the square root of 400.. It is 20. That's correct right?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (jdoe0001):

 \(\bf c^2=a^2+{\color{blue}{ b}}^2\implies c^2-a^2={\color{blue}{ b}}^2\implies \sqrt{c^2-a^2}={\color{blue}{ b}}\to \sqrt{400}\to 20\) yes

OpenStudy (solomonzelman):

yes 20 is correct.

OpenStudy (solomonzelman):

Another way to know this, 15 b 25 3 ? 5 a multiply of 3,4,5 :)

OpenStudy (solomonzelman):

* mutliple